# Jacobian-Enhanced Neural Networks

### v2

@@ Abstract
Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. Their main benefit is better accuracy with fewer training points compared to standard neural networks. These attributes are particularly desirable in the field of computer-aided design, where there is often the need to replace computationally expensive, physics-based models with fast running approximations, known as surrogate models or meta-models. Since a surrogate emulates the original model accurately in real time, it yields a computational speed benefit that can be used to carry out orders of magnitude more function calls quickly. However, in the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which is a critical property for one important use-case: surrogate-based optimization. This work derives the complete theory and exemplifies its superiority over standard neural nets for surrogate-based optimization.
Index Terms:
Jacobian, neural network, gradient-enhanced, multi-layer perceptron, surrogate-based optimization, surrogate model, meta-model.Nomenclature
@@
I Introduction
@@
Jacobian-Enhanced Neural Networks (JENN) are fully connected multi-layer perceptrons, whose training process is modified to predict partial derivatives accurately. This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of a gradient for vector functions with more than one output. The main benefit of jacobian-enhancement (synonymous to gradient-enhancement) is better accuracy with fewer training points compared to standard fully connected neural nets. This paper develops the theory behind JENN and quantifies its benefits for the intended use-case. It complements previous work in the following ways. First, it provides a concise but complete derivation, including equations in vectorized form (see appendix), without relying on automatic differentiation. Second, it provides a deep-learning formulation that can handle any number of inputs, hidden layers, or outputs. Finally, through judicious use of hyperparameters, it enables weighting; i.e. a process by which individual sample points can be prioritized differently in situations that call for it. The latter, in particular, gives rise to the concept of polishing (described in the application section) which offers an additional tuning mechanism to further improve Surrogate-Based Optimization (SBO) in gradient-descent applications.
I-A Intended Use-Case
@@
JENN is intended for the field of computer aided design, where there is often a need to replace computationally expensive, physics-based models with fast running approximations, commonly referred to as surrogate models or meta-models. Since a surrogate emulates the original model accurately in fractions of second, it yields a speed benefit that can be used to carry out orders of magnitude more function calls quickly. This allows designers to explore a broader tradespace, and more readily quantify the impact of various parameters in order to better understand the performance of the system under design. While this is true of any surrogate modeling approach, JENN has an additional benefit that most surrogate modeling approaches lack: the ability to predict partial derivatives with improved accuracy. This is a critical property for one important use-case: surrogate-based optimization. The field of engineering is rich in design applications of such a use-case with entire book chapters dedicated to it [1, 2, 3, 4], review papers [5, 6, 7, 8], and countless examples to be found, particularly in the field of aerospace engineering [9, 10, 11, 12].
I-B Limitations
@@
First and foremost, gradient-enhanced methods require factors and responses to be smooth and continuous. Second, the value proposition of a surrogate is that the computational expense of generating training data to fit the model is much less than the computational expense of directly performing the analysis with the original model itself, which is not always true, especially when computing partials. Gradient-enhanced methods are therefore only beneficial if the cost of obtaining partials is not excessive in the first place (e.g. adjoint methods [13] in computational fluid dynamics), or if the need for accuracy outweighs the cost of computing the partials. One should therefore carefully weigh the benefit of gradient-enhanced methods relative to the needs of their application.
I-C Supporting Software
An open-source Python library that implements the theory can be found at https://pypi.org/project/jenn/. There exist many excellent deep learning frameworks, such as TensorFlow [14] and PyTorch [15], which are more performant and versatile than JENN, as well as many all-purpose machine learning libraries, such as scikit-learn [16]. However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, which requires non-trivial effort to be implemented in those tools. The present library seeks to close that gap. It is intended for engineers with limited time, who have a need to predict partials accurately, and who are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry111Note that GENN from the Surrogate Modeling Toolbox [17] is JENN with a different API (present author contribution).
II Previous Work
@@
A cursory literature research reveals that the field of gradient-enhanced regression is predominantly focused on Gaussian processes. Dedicated book chapters can be found about gradient-enhanced Kriging models [2], salient open-source libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods [17], and even review papers about gradient-enhanced meta-models only mention neural networks in passing [18]. However, early work by Sellar and Batill [19] can be found in the late 1990’s, which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization. They conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all work surveyed for this paper.
@@
Follow-on research by that Liu and Batill [20] provides the mathematical derivation of the equations for neural net gradient-enhancement, but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility. This shortcoming is overcome by Giannakoglou et al. [10] in the late 2000’s who developed a deep neural-network formulation with gradient-enhancement and use it for aerodynamic design applications. All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
@@
Recent work by Bouhlel et al. [11] in 2020 takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic-differentiation tools. Bouhlel et al. successfully apply gradient-enhanced neural networks to aerodynamic use-cases and re-affirm the previous conclusion that such techniques improve accuracy and reduce the amount of training data needed, provided the cost of obtaining the gradient is not prohibitive in the first place. While deep learning frameworks have obvious advantages (e.g. parallel computing), there is complementary research value in developing the equations for back-propagation, which is a special case of automatic differentiation for scalar functions. Stand-alone equations are reproducible without third-party tools and convey deeper insight into the mechanics behind an algorithm, which others can build upon.
@@
The works surveyed so far have focused on prediction accuracy, but Hoffman et al. [21] take a different approach by focusing on robustness instead. They introduce the concept of Jacobian regularization for classification, where the key idea is to ensure robustness of models against input perturbations due to adversarial attacks; for example, a purposefully modified input image to avoid detection. However, robustness is outside the scope of the present work, which is to predict design partials accurately given high quality inputs.
@@
Finally, the most recent work on gradient-enhanced neural nets focuses on gradient-enhanced Physics-Informed Neural Networks [22] (PINN). This novel research uses the residuals (and its partials) of Partial Differential Equations (PDE), governing some physical system of interest, as the loss function for training neural nets (instead of using the classic least squares estimator). However, PINN is outside the scope of the present work, which is to develop a general purpose algorithm to fit any smooth function, regardless of what it represents, even if it is not physics-informed.
III Mathematical Derivation
@@
Jacobian-Enhanced Neural Networks (JENN) follow the same steps as standard Neural Networks (NN), shown in Fig. 1, except that the cost function used during parameter update is modified to account for partial derivatives, with the consequence that back propagation must be modified accordingly. This section provides the mathematical derivation for Jacobian-enhancement.
III-A Notation
@@
For clarity, let us adopt a similar notation to the one used by Andrew Ng [23] where superscript denotes the layer number, is the training example, and subscript indicates the node number in a layer. For instance, the activation of the node in layer evaluated at example would be denoted:
@@
Derivatives with respect to input are denoted by the prime symbol and subscript . For example:
@@
III-B Normalize Training Data
Normalizing the training data can greatly improve performance when either inputs or outputs differ by orders of magnitude. A simple way to improve numerical conditioning is to subtract the mean and divide by the standard deviation of the data. The normalized data then becomes:
This approach works well, except when all values are very close such that the denominator vanishes as standard deviation goes to zero. In that case, alternative normalization methods (or none) should be considered.
III-C Forward Propagation
@@
III-D Parameter Update
The neural network parameters are learned by solving the following, unconstrained optimization problem:
where it should be understood that is short hand for all parameters of the neural network. The cost function is given by the following expression, where is a hyper-parameter to be tuned which controls regularization:
@@ III-D Parameter Update
where the hyper-parameter is the learning rate. In practice, some improved version of gradient-descent would be used instead, such as ADAM [24], but the former is easier to explain without loss of generality. The loss function is taken to be a modified version of the Least rquares Estimator (LSE), augmented with an extra term to account for partial derivatives:
@@
Note that two more hyper-parameters have been introduced: and . They act as weights that can be used to prioritize subsets of the training data in situations that call for it. By default, , meaning that all data is prioritized equally, which is the common use-case. However, in some situations, there may be an incentive to prioritize certain examples over others, as will be shown in the results section. This formulation offers the flexibility to do so.
III-E Backward Propagation
@@
Forward propagation seeks to compute the gradient of the predicted outputs with respect to (abbreviated w.s.t.) the inputs . By contrast, back-propagation computes the gradient of the cost function with respect to the neural net parameters :
III-E1 Loss Function Derivatives w.s.t. Neural Net Parameters
@@
In a first step, assume the quantities and are known and consider the gradient of the loss function with respect to the parameters from layer . Applying the chain rule to Eq. III-D yields:
@@
where the superscript has been dropped for notational convenience, with the understanding that all equations to follow apply to example . The various partial derivatives in Eq. 11 follow from Eq. 3:
rubstituting these expressions into Eq. 11 yields:
III-E2 Loss Function Derivatives w.s.t. to Neural Net Activations
@@
All quantities in Eq. 15 are known, except for and which must be obtained recursively by propagating information backwards from one layer to the next, starting with the output layer . From Eq. III-D, it follows that:
@@
The derivatives for the next layer are obtained through the chain rule. According to Eq. III-D, the loss function depends on activations and associated partials , which are themselves functions of and . Hence, according to the chain rule, current layer derivatives w.s.t. to previous layer activations are:
@@
Further applying the chain rule to the intermediate quantities and using Eq. 3, one finds that:
@@
All necessary equations to compute the derivatives of the loss function w.s.t. parameters in any layer are now known. Derivatives are obtained by working our way back recursively across layers, repeatedly applying Eq. 17 after passing the solution in the current layer to the preceding layer as follows:
III-E3 Cost Function Derivatives w.s.t. Parameters
IV Verification and Validation
The theory developed in the previous section was implemented into an installable Python library for which a link was provided earlier. This section shows validation results333code to reproduce available at https://codeocean.com/capsule/9360536/tree for selected test functions in Fig. 2, where training data is shown as dots. These validation results compare JENN against a standard Neural Net (NN), defined as the equivalent neural network without gradient-enhancement.
@@ IV Verification and Validation
All else being equal, it can clearly be seen that JENN outperforms NN in all cases, re-affirming the general concensus that gradient-enhancement achieves better accuracy with less data, so long as the computational cost of obtaining the Jacobian is not prohibitive in the first place. In Figs. 2a and 2b, where there are only three and four training samples respectively, NN fails to even capture the correct trend whereas JENN predicts the response almost perfectly. In Fig. 2d, where there are 100 samples, NN correctly captures the trend but contour lines are distorted, implying more samples would be needed to improve prediction. By contrast, under the same conditions, JENN predicts the Rastrigin function [25] almost perfectly.
@@
Finally, it was also verified that runtime scales as , as shown in Fig. 2c. This plot was obtained using the Rastrigin function to generate progressively larger datasets, train a new model, and record runtime on a MacBook Pro (2021).
V Example Application: Surrogate-Based Optimization

### v3

@@ Abstract
Jacobian-Enhanced Neural Networks (JENN) are densely connected multi-layer perceptrons, whose training process is modified to predict partial derivatives more accurately. Their main benefit is better accuracy with fewer training points compared to standard neural networks. These attributes are particularly desirable in the field of design optimization, where there is often the need to replace computationally expensive models with fast running approximations, known as surrogate models (or metamodels), to speed up solutions. In the special case of gradient-enhanced methods, there is the additional value proposition that partial derivatives are accurate, which yields improved accuracy for gradient-based optimization. This work develops the theory of Jacobian-Enhanced Neural Nets and applies it to an airfoil optimization problem in order to illustrate its benefits over the status quo.
keywords:
Jacobian , gradient-enhanced , neural networks , surrogate model , aerodynamic shape optimization[ELSYS]organization=Sr Researcher Engineer at Georgia Tech Research Institute, city=Atlanta, state=GA, postcode=30332, country=USA
A neural network formulation capable of predicting partial derivatives more accurately with significantly fewer training data samples
Practical illustration through applied aerospace engineering example, taken to be aerodynamic shape optimization using CFD
Complete mathematical derivation of the underlying theory for arbitrary deep network, including vectorized equations for computer programming, with open-source Python code implementation
Quantification of advantages & limitations through canonical example, including analysis of noisy partials and performance on benchmark optimization problem
Convenient framework to handle missing partials, believed to be a common use-case of practical importance.
Nomenclature
@@
1 Introduction
Aerodynamic shape optimization is concerned with maximizing aerodynamic performance by changing the shape of a body subjected to specific flow conditions. The state of the art uses computational fluid dynamics and adjoint methods to achieve this goal (Jameson1988; Kenway2019). These methods scale very well, enabling optimization problems with hundreds of decision variables to be solved as efficiently as problems with a handful. However, the cost of individual function calls remains an impediment: a single flow solution can take several hours depending on the use case. For this reason, there has been significant interest in accelerating numerical computations through machine learning in aerodynamic shape optimization (Li2022), and the present work focuses on one method in particular: Jacobian-Enhanced Neural Networks (JENN).
1.1 Overview
@@
JENN are fully connected multi-layer perceptrons, whose training process has been modified to predict partial derivatives more accurately. This is accomplished by minimizing a modified version of the Least Squares Estimator (LSE) that accounts for Jacobian prediction error, where a Jacobian extends the concept of the gradient for functions of more than one output. The main benefit of Jacobian enhancement is better accuracy with fewer training points compared to standard fully connected neural nets. This paper develops the theory behind JENN and applies it to an airfoil optimization problem in order to illustrate its benefits over status quo neural networks. It complements previous work on gradient-enhanced neural nets in three ways. First, it provides a concise and complete derivation of the theory, including equations in vectorized form, without relying on automatic differentiation. Second, it provides a deep-learning formulation that can handle any number of inputs, hidden layers, or outputs. Third, through judicious use of hyperparameters, it enables individual sample points to be prioritized differently in situations that call for it, the utility of which is illustrated in the application section.
1.2 Intended Use-Case
@@
Even though JENN was developed with aerodynamic shape optimization in mind, it is a general method that applies to any engineering problem for which partial derivatives are defined. It is intended for the field of computer aided design, where there is often a need to replace computationally expensive simulation models with fast running approximations, commonly referred to as surrogate models or metamodels. Since a surrogate emulates the original model with close approximation in a fraction of a second, it yields a speed benefit that enables orders of magnitude more function calls to be carried out quickly. This allows designers to explore a tradespace more fully, in order to better understand the impact of design variables on the performance of the system under design. While this is true of any surrogate modeling approach, JENN has an additional benefit that most surrogate modeling approaches lack: the ability to predict partial derivatives more accurately. This is a desirable property for one use case in particular: surrogate-based optimization. The field of engineering is rich in design applications of such a use case with entire book chapters dedicated to it (Keane2005; Forrester2011; Iuliano2016; Martins_Ning_2021), review papers (Queipo2005; haridy2023review; su14073867; Wang2006), and countless examples to be found, particularly in the field of aerospace engineering (Robinson2008; Giannakoglou2006; Bouhlel2020; Nagawkar2022).
1.3 Limitations
@@
First and foremost, gradient-enhanced methods require factors and responses to be smooth and continuous. Second, the value proposition of a surrogate is that the computational expense of generating training data to fit the model is much less than the computational expense of directly performing the analysis with the original model itself, which is not always true, especially when computing partials. Gradient-enhanced methods are therefore only beneficial if the cost of obtaining partials is not excessive in the first place or if the need for accuracy outweighs the cost of computing the partials; see Martins et al. (Martins2013a) for an excellent review of all existing methods for computing the derivatives of computational models.
2 Previous Work
@@
A cursory literature research reveals that the field of gradient-enhanced learning is predominantly focused on Gaussian processes. Dedicated book chapters can be found about gradient-enhanced Kriging models (Forrester2011), salient open-source libraries have been almost exclusively devoted to gradient-enhanced Bayesian methods (saves2024smt), and even review papers about gradient-enhanced metamodels only mention neural networks in passing (Laurent2019). However, early work by (Sellar1996) can be found which uses gradient-enhanced neural network approximations in the context of Concurrent Subspace Optimization (CSSO), a type of optimization architecture in the field of multi-disciplinary design optimization. The authors of that work conclude that gradient-enhanced neural networks are more accurate and require significantly less training data than their non-enhanced counterparts, which seems to be the general consensus in all surveyed work to follow.
@@
Follow-on research by (Liu2000) further provides the mathematical derivation of the equations for neural network gradient-enhancement, but the formulation is not general; it is limited to shallow neural networks with only two layers, which is of limited utility. This shortcoming is overcome by (Giannakoglou2006) who developed a deep neural network formulation with gradient enhancement and used it for aerodynamic design applications. All necessary equations to reproduce their algorithm are provided, but their formulation is limited to only one output and the vectorized form is missing for computer programming.
@@
More recent work by (Bouhlel2020) takes a different approach, relying instead on established deep-learning frameworks to delegate back-propagation to third-party, automatic differentiation tools. Using TensorFlow (tensorflow2015), gradient-enhanced neural networks are successfully used to accelerate airfoil shape optimization, culminating into webfoil111https://webfoil.engin.umich.edu/; a useful web App for fast airfoil shaping. Overall, those authors reaffirm the previous conclusion that gradient-enhanced techniques improve accuracy and reduce the amount of training data required for good fits. While deep learning frameworks have obvious advantages (e.g. parallel computing), there is complementary research value in developing the equations for back-propagation from scratch, which is one of the objectives of this paper. Stand-alone equations are reproducible without third-party tools and convey helpful insight into the mathematics behind a specific algorithm. Backprop is a special case of automatic differentiation for scalar functions.
@@
The works surveyed so far have focused on prediction accuracy, but a keyword search for Jacobian and neural networks also points to the concept of Jacobian regularization (hoffman2019robust) for classification. The key idea is to ensure robustness of models against input perturbations due to adversarial attacks (e.g. a purposefully modified input image to avoid detection) by using Jacobian information to reduce sensitivity. However, robustness is outside the scope of the present work, which seeks to predict partials accurately for design applications given high quality inputs.
@@
Finally, the most recent work on gradient-enhanced neural networks focuses on gradient-enhanced Physics-Informed Neural Networks (PINN) (yu2022), which uses the residuals of partial differential equations as the loss function for training neural nets. It is a deep learning framework for predicting systems governed by nonlinear partial differential equations, in which the neural network effectively learns to solve a system of equations rather than merely emulating a dataset. However, PINN falls outside the scope of the present work, which seeks to develop a general purpose algorithm to fit any smooth function regardless of what it represents, be it physics-informed or not.
3 Mathematical Derivation
@@
Jacobian-Enhanced Neural Networks (JENN) follow the same steps as standard Neural Networks (NN), shown in Fig. 1, except that the cost function used during parameter update is modified to account for partial derivatives, with the consequence that back propagation must be modified accordingly. This section provides the mathematical derivation.
3.1 Notation
@@
For clarity, let us adopt a similar notation to the one used by (Ng2017) where superscript denotes the layer number, is the training example, and subscript indicates the node number in a layer. For instance, the activation of the node in layer evaluated at example would be denoted:
@@
Derivatives w.r.t.input are denoted by the prime symbol and subscript , e.g.
@@
3.2 Normalize Training Data
Normalizing the training data can greatly improve performance when either inputs or outputs differ by orders of magnitude. A simple way to improve numerical conditioning is to subtract the mean and divide by the standard deviation of the data, provided the denominator is not close to zero. The normalized data then becomes:
3.3 Forward Propagation
@@
3.4 Parameter Update
Neural network parameters are learned by solving an unconstrained optimization problem:
where it should be understood that is short hand for all parameters of the neural network. The cost function is given by the following expression, where is a hyperparameter to be tuned which controls regularization:
@@ 3.4 Parameter Update
where the hyperparameter is the learning rate. In practice, some improved version of gradient-descent would be used instead, such as ADAM (kingma2017adam), but the former is easier to explain without loss of generality. The loss function is taken to be a modified version of the Least Squares Estimator (LSE), augmented with an extra term to account for partial derivatives:
@@
Note that two more hyperparameters have been introduced: and . They act as weights that can be used to prioritize subsets of the training data in situations that call for it (example later). By default, , meaning that all data is prioritized equally.
3.5 Backward Propagation
@@
Forward propagation seeks to predict the gradient of outputs with respect to (abbreviated w.r.t.) inputs . By contrast, back-propagation computes the gradient of the cost function w.r.t. neural net parameters :
3.5.1 Loss Function Derivatives w.r.t. Neural Net Parameters
@@
In a first step, assume the quantities and are known and consider the gradient of the loss function w.r.t.the parameters from layer . Applying the chain rule to Eq. 9 yields:
@@
where the superscript has been dropped for notational convenience, with the understanding that all equations to follow now apply to example . The various partial derivatives in Eq. 11 follow from Eq. 3:
Substituting these expressions into Eq. 11 yields:
3.5.2 Loss Function Derivatives w.r.t. to Neural Net Activations
@@
All quantities in Eq. 15 are known, except for and which must be obtained recursively by propagating information backwards from one layer to the next, starting with output layer . From Eq. 9, it follows that:
@@
The derivatives for the next layer are obtained through the chain rule. According to Eq. 9, the loss function depends on activations and associated partials , which are themselves functions of and . Hence, according to the chain rule, previous layer derivatives w.r.t. to current layer activations are:
@@
Further applying the chain rule to the intermediate quantities and using Eq. 3, one finds:
@@
All necessary equations to compute the derivatives of the loss function w.r.t. parameters in any layer are now known. Derivatives are obtained by working our way back recursively across layers, repeatedly applying Eq. 17 after passing the solution in the current layer to the preceding layer as follows:
3.5.3 Cost Function Derivatives w.r.t. Parameters
Putting it all together, the derivative of the cost function w.r.t. neural network parameters are obtained by substituting Eq. 15 into Eq. 10, where the subscripts and are referenced to layer :
A summary of key equations is available in A in vectorized form for computer programming, which have already been implemented in an open-source library called JENN developed for this work.
4 Implementation
An open-source Python library that implements JENN can be found online333https://pypi.org/project/jenn/. There exist many excellent deep learning frameworks, such as TensorFlow (tensorflow2015) and PyTorch (PyTorch), as well as many all-purpose machine learning libraries such as scikit-learn (sklearn2011). However, gradient-enhancement is not inherently part of those popular frameworks, let alone the availability of prepared functions to predict partials after training is complete, which requires non-trivial effort to be implemented in those tools. The present library seeks to close that gap. It is intended for non-practitioners, who have access to partials and are seeking a minimal-effort Application Programming Interface (API) with low-barrier to entry to take advantage of them. JENN focuses exclusively on neural networks and serves as an upstream dependency to the popular Surrogate Modeling Toolbox (SMT) (saves2024smt) specializing in gradient-enhanced methods in general.
4.1 Verification and Validation
Using canonical test functions, Fig. 2 compares JENN against a standard Neural Net (NN), taken to be the same neural network with gradient-enhancement turned off (i.e. ). All else being equal, it can be seen that JENN outperforms NN in all cases, reaffirming the general concensus that gradient-enhancement achieves better accuracy with less data.
@@ 4.1 Verification and Validation
In Figs. 2a and 2b, where there are only three and four training samples respectively, NN fails to even capture the correct trend whereas JENN predicts the response almost perfectly. In Fig. 2c, where there are 100 samples, NN correctly captures the trend but contour lines are distorted, implying more samples would be needed to improve prediction. By contrast, under the same conditions, JENN predicts the Rastrigin function (rastrigin1974systems) almost perfectly. The code to reproduce all results in this section is available online444code to reproduce available at https://codeocean.com/capsule/9360536/tree.
4.2 Run Time
@@
It was empirically verified that runtime scales linearly as with sample size, as shown in Fig. 3. This plot was obtained using the Rastrigin function (rastrigin1974systems) to generate progressively larger datasets, train a new model, and record runtime on a MacBook Pro (2021).
4.3 Surrogate-Based Optimization